### v1

# FreeSplat: Generalizable 3D Gaussian Splatting Towards Free-View Synthesis of Indoor Scenes

Abstract
Empowering 3D Gaussian Splatting with generalization ability is appealing. However, existing generalizable 3D Gaussian Splatting methods are largely confined to narrow-range interpolation between stereo images due to their heavy backbones, thus lacking the ability to accurately localize 3D Gaussian and support free-view synthesis across wide view range. In this paper, we present a novel framework FreeSplat that is capable of reconstructing geometrically consistent 3D scenes from long sequence input towards free-view synthesis. Specifically, we firstly introduce Low-cost Cross-View Aggregation achieved by constructing adaptive cost volumes among nearby views and aggregating features using a multi-scale structure. Subsequently, we present the Pixel-wise Triplet Fusion to eliminate redundancy of 3D Gaussians in overlapping view regions and to aggregate features observed across multiple views. Additionally, we propose a simple but effective free-view training strategy that ensures robust view synthesis across broader view range regardless of the number of views. Our empirical results demonstrate state-of-the-art novel view synthesis peformances in both novel view rendered color maps quality and depth maps accuracy across different numbers of input views. We also show that FreeSplat performs inference more efficiently and can effectively reduce redundant Gaussians, offering the possibility of feed-forward large scene reconstruction without depth priors. Our code will be made open-source upon paper acceptance.
Figure 1: Comparison between FreeSplat and previous methods. pixelSplat [1] and MVSplat [2] fail to reconstruct geometrically consistent global 3D Gaussians, while our FreeSplat is proposed to accurately localize 3D Gaussians from long sequence input and support free view synthesis.
1 Introduction
Recent advancements has emerged [3, 4, 5, 6] in reconstructing 3D scenes from multiple viewpoints. Based on ray-marching-based volume rendering, Neural Radiance Fields [3, 7, 8, 9] is capable of learning the implicit 3D geometry and radiance fields without depth information. Nonetheless, computational cost remains to be the inherent bottleneck in ray-marching-based volume rendering, preventing it from real-time rendering. 3D Gaussian Splatting [10, 11, 12, 13] has recently been proposed as an efficient representation for photorealistic reconstruction of 3D scenes from multi-views. The explicit representation of 3D Gaussians are optimized to be densified in the textured regions, and the rasterization-based volume rendering avoids the costly ray marching scheme. Consequently, 3D Gaussian Splatting has achieved real-time rendering of high-quality images from novel views. Nonetheless, vanilla 3D Gaussian Splatting lacks generalizability and requires per-scene optimization.
Several attempts [1, 2, 14, 15, 16, 17] have been made to give 3D Gaussian Splatting generalization ability. Despite showing promising performance, these methods are limited to narrow-range scene-level view interpolation [1, 2, 15] and object-centric synthesis [14, 16]. The primary reason for the limitation is that these existing methods depend on dense view matching across multi-view images with transformers to predict Gaussian primitives, which consequently becomes computationally intractable with longer sequences and thus restricting the supervision of these methods to narrow-range interpolated views. As we show in Figure 3, supervision by narrow-range interpolated views often result in poorly localized 3D Gaussians that can become floaters when rendered from extrapolated views. Additionally, the problem is further aggrevated by existing methods typically merging multi-view 3D Gaussians through simple concatenation and thus inevitably lead to noticeable redundancy in overlapping areas (cf. Table 2). In view of the above-mentioned problems, it is therefore imperative to design a method that is capable of long sequence reconstruction of global 3D Gaussians, which has the significant potential of supporting real-time rendering from arbitrary poses.
In this paper, we propose FreeSplat tailored for indoor long sequence free view synthesis. Unlike existing methods limited to view interpolation in narrow ranges, our method can effectively reconstruct explicit global 3DGS for novel view synthesis across wide view ranges. Our pipline consists of Low-cost Cross-View Aggregation and Pixel-wise Triplet Fusion (PTF). In Low-cost Cross-View Aggregation, we introduce efficient CNN-based backbones and adaptive cost volumes formulation among nearby views for low-cost feature extraction and matching, then we leverage a Multi-Scale Feature Aggregation structure to broaden the receptive field of cost volume and predict Depths and Gaussian Triplets. Subsequently, we present Pixel-wise Alignment with progressive Gaussian fusion in PTF to adaptively fuse local Gaussian Triplets from multi-views and avoid Gaussian redundancy in the overlapping regions. Moreover, due to our efficient feature extraction and matching, we propose a Free-View Training (FVT) strategy to disentangle generalizable 3DGS with specific number of views and train the model on long sequences.
The contributions of our paper are summarized as follows:
  1. 1.
    We present Low-cost Cross-View Aggregation to predict initial Gaussian triplets, where the low computational cost makes it possible for feature matching between more nearby views and training on long sequence reconstruction;
  2. 2.
    We propose Pixel-wise Triplet Fusion to fuse Gaussian triplets, which can effectively reduce the Gaussian redundancy in the overlapping regions and aggregate multi-view 3D Gaussian latent features;
  3. 3.
    To the best of our knowledge, we are the first to explore generalizable 3DGS for long sequence reconstruction. Extensive experiments on indoor dataset ScanNet [18] and Replica [19] demonstrate our superiority on both image rendering quality and novel view depth rendering accuracy when given different lengths of input views.
2 Related Work
Novel View Synthesis. Traditional attempts in novel view synthesis mainly employed voxel grids [20, 21] or multiplane images [22]. Recently, Neural Radiance Fields (NeRF) [3, 5, 23, 24, 25] have drawn growing interest using ray-marching-based volume rendering to backpropagate image color error to the implicit geometry and radiance fields, such that the 3D geometry can be implicitly learned to satisfy the multi-view color consistency. Nonetheless, one inherent bottleneck of NeRFs-based method is the computation intensity of ray marching, which requires the costly volume sampling in the implicit fields for each pixel during rendering. To this end, recently 3DGS [10, 26, 27, 11] have attracted increasing attention due to its high efficiency and photorealistic rendering. Instead of relying on MLPs to represent the coordincate-based implicit fields, 3DGS learns an explicit field using a set of 3D Gaussians. They optimize the 3D Gaussians parameters and perform adaptive densify control to fit to the given set of images, such that the 3D Gaussians are encouraged to perform densification only in the textured regions and refrain from over-densification. During rendering, 3DGS performs tile-based rasterization to differentiably accumulate color images from the explicit 3D Gaussian primitives, which is significantly faster than the ray-marching-based volume rendering and achieves real-time rendering speed.
Generalizable Novel View Synthesis. Another drawback of the traditional NeRF-based and 3DGS-based methods is the requirement of per-scene optimization instead of direct feeding-forward. To this end, there have been a line of work [28, 8, 7, 29] focusing on learning effective priors to predict 3D geometry from given images in a feed forward fashion, where the common practice is to project ray-marching sampled points onto given source views to aggregate multi-view features, conditioning the prediction of the implicit fields on source views instead of point coordinates. Recently, there have also been attempts towards generalizable 3DGS [1, 17, 2, 15, 14]. pixelSplat [1] and GPS-Gaussian [17] propose to predict pixel-aligned 3D Gaussian parameters in feed forward fashion. MVSplat [2] replaces the epipolar line transformer of pixelSplat with a lightweight cost volume to perform more efficient image encoding. GGRt [15] concatenates pixelSplat predicted 3D Gaussians in a sequence of images and simultaneously perform pose optimization. latentSplat [14] encodes 3D Variational Gaussians and leverages a discriminator to help produce more indistinguable images. Nonetheless, existing methods do not reconstruct the global 3D Gaussians from arbitrary length of inputs, and are limited to view interpolation [1, 2, 17, 15] or object/human-centric scenes [17, 14]. In contrary, in this paper we focus on reconstructing large scenes from arbitrary length of inputs without depth priors, unleashing the potential of generalizable 3DGS for large scene explicit representation.
Indoor Scene Reconstruction. One line of efforts in feed-forward indoor scene reconstruction focuses on extracting 3D mesh using voxel volumes [30, 31, 32] and TSDF-fusion [33], while do not perform photorealistic novel view synthesis. On the other hand, the SLAM-based methods [34, 35, 36] require dense sequence of RGB-D input and per-scene tracking and mapping. Another paradigm of 3D reconstruction [37, 38, 39] learns implicit Signed Distance Fields from RGB input, while demanding intensive per-scene optimization. Another recent work SurfelNeRF [40] learns a feed-forward framework to map a sequence of images to 3D surfels which support photorealistic image rendering, while they rely on external depth estimator or ground truth depth maps. In contrary, we propose an end-to-end model without ground truth depth map input or supervision, enabling accurate 3D Gaussian localization using only photometric losses.
3 Preliminary
Vanilla 3DGS. 3D-GS [10] explicitly represents a 3D scene with a set of Gaussian primitives which are parameterized via a 3D covariance matrix $\bm{\Sigma}$ and mean $\bm{\mu}$:
$G(\textbf{p})=\mathrm{exp}(-\frac{1}{2}\left(\bm{p}-\bm{\mu}\right)^{\top}\bm{\Sigma}^{-1}\left(\bm{p}-\bm{\mu}\right)),$ (1)
where $\bm{\Sigma}$ is decomposed into $\bm{\Sigma}=\bm{\mathrm{R}}\bm{\mathrm{S}}\bm{\mathrm{S}}^{\top}\bm{\mathrm{R}}^{\top}$ using a scaling matrix $\bm{\mathrm{S}}$ and a rotation matrix $\bm{\mathrm{R}}$ to maintain positive semi-definiteness. During rendering, the 3D Gaussian is transformed into the image coordinates with world-to-camera transform matrix $\bm{\mathrm{W}}$ and projected onto image plane with projection matrix $\bm{\mathrm{J}}$, and the 2D covariance matrix $\bm{\mathrm{\Sigma}}^{\prime}$ is computed as $\bm{\Sigma}^{\prime}=\bm{\mathrm{J}}\bm{\mathrm{W}}\bm{\mathrm{\Sigma}}\bm{\mathrm{W}}^{\top}\bm{\mathrm{J}}^{\top}$. We then obtain a 2D Gaussian $G^{2D}$ with the covariance $\bm{\Sigma}^{\prime}$ in 2D, and the color rendering is computed using point-based alpha-blending on each ray:
$\bm{\mathrm{C}}(\bm{\mathrm{x}})=\sum_{i\in N}\bm{\mathrm{c}}_{i}\alpha_{i}G^{2D}_{i}(\bm{\mathrm{x}})\prod_{j=1}^{i-1}(1-\alpha_{j}G^{2D}_{j}(\bm{\mathrm{x}})),$ (2)
where $N$ is the number of Gaussian primitives, $\alpha_{i}$ is a learnable opacity, and $\bm{\mathrm{c}}_{i}$ is view-dependent color defined by spherical harmonics (SH) coefficients $\bm{\mathrm{s}}$. The Gaussian parameters are optimized by a photometric loss to minimize the difference between renderings and image observations.
Generalizable 3DGS. Unlike vanilla 3DGS that optimizes per-scene Gaussian primitives, recent generalizable 3DGS [1, 17] predict pixel-aligned Gaussian primitives $\{\bm{\Sigma},\bm{\alpha},\bm{\mathrm{s}}\}$ and depths $\bm{d}$, such that the pixel-aligned Gaussian primitives can be unprojected to 3D coordinates $\bm{\mu}$. The Gaussian parameters are predicted by 2D encoders, which are optimized by the photometric loss through rendering from novel views. However, existing methods are still limited to view interpolation within narrow view range, which leads to inaccurately localized 3D Gaussians that fail to support large scene reconstruction and view extrapolation (cf. Figure 1, 3). To this end, we propose FreeSplat towards global 3D Gaussians reconstruction with accurate localization that supports free-view synthesis.
4 Our Methodology
Figure 2: Framework of FreeSplat. Given input sparse sequence of images, we construct cost volumes between nearby views and predict depth maps and corresponding feature maps, followed by unprojection to Gaussian triplets with 3D positions. We then propose Pixel-aligned Triplet Fusion (PTF) module, where we progressively aggregate and update local/global Gaussian triplets based on pixel-wise alignment. The global Gaussian triplets can be later decoded into Gaussian parameters.
4.1 Overview
The overview of our method is illustrated in Figure 2. Given a sparse sequence of RGB images, we build cost volumes adaptively between nearby views, and predict depth maps to unproject the 2D feature maps into 3D Gaussian triplets. We then propose the Pixel-aligned Triplet Fusion (PTF) module to progressively align the global with the local Gaussian triplets, such that we can fuse the redundant 3D Gaussians in the latent feature space and aggregate cross-view Gaussian features before decoding. Our method is capable of efficiently exchanging cross-view features through cost volumes, and progressively aggregating per-view 3D Gaussians with cross-view alignment and adaptive fusion.
4.2 Low-cost Cross-View Aggregation
Efficient 2D Feature Extraction. Given a sparse sequence of posed images $\{\bm{I}^{t}\}_{t=1}^{T}$, we first feed them into a shared 2D backbone to extract multi-scale embeddings $\bm{F}_{e}^{t}$. To decrease and computational cost in plane sweeping, we train a matching net [41] to embed the low-dimensional matching feature $\bm{F}_{m}^{t}$. Unlike [1, 2] which rely on patch-wise transformer-based backbones [42, 43] that can lead to quadratically expensive computations, we leverage pure CNN-based backbones [44, 41] for 2D feature extraction for efficient performance on higher resolution inputs.
Adaptive Cost Volume Formulation. To explicitly integrate camera pose information given arbitrary length of input images, we propose to adaptively build cost volumes between nearby views. For current view $\bm{I}^{t}$ with pose $\bm{P}^{t}$ and matching feature $\bm{F}_{m}^{t}\in\mathbb{R}^{C_{m}\times\frac{H}{4}\times\frac{W}{4}}$, we adaptively select its $N$ nearby views $\{\bm{I}^{t_{n}}\}_{n=1}^{N}$ with poses $\{\bm{P}^{t_{n}}\}_{n=1}^{N}$ based on pose proximity, and construct cost volume via plane sweep stereo [45, 46]. Specifically, we define a set of $K$ virtual depth planes $\{d_{k}\}_{k=1}^{K}$ that are uniformly spaced within $[d_{near},d_{far}]$, and warp the nearby view features to each depth plane $d_{k}$ of current view:
$\tilde{\bm{F}}_{m}^{t_{n},k}=\mathrm{Trans}(\textbf{P}^{t_{n}},\textbf{P}^{t})\bm{F}_{m}^{t_{n}},$ (3)
where $\mathrm{Trans}(\textbf{P}^{t_{n}},\textbf{P}^{t})$ is the transformation matrix from view $t_{n}$ to $t$. The cost volume $\bm{F}_{\mathrm{cv}}^{t}\in\mathbb{R}^{K\times\frac{H}{4}\times\frac{W}{4}}$ is then defined as:
$\bm{F}_{\rm{cv}}^{t}(k)=f_{\theta}\left((\frac{1}{N}\sum_{n=1}^{N}\mathrm{cos}(\bm{F}_{m}^{t},\tilde{\bm{F}}_{m}^{t_{n},k}))\oplus(\frac{1}{N}\sum_{n=1}^{N}\tilde{\bm{F}}_{m}^{t_{n},k})\right),$ (4)
where $\bm{F}_{\rm{cv}}^{t}[k]$ is the $k$-th dimension of $\bm{F}_{\rm{cv}}^{t}$, $\mathrm{cos}(\cdot)$ is the cosine similarity, $\oplus$ is feature-wise concatenation, and $f_{\theta}(\cdot)$ is a $1\times 1\;\mathrm{CNN}$ mapping to dimension of $1$.
Multi-Scale Feature Aggregation. The embedding of the cost volume plays a significant part to accurately localize the 3D Gaussians (cf. Table 5, 6). To this end, inspired by previous depth estimation methods [47, 33], we design an multi-scale encoder-decoder structure, such that to fuse multi-scale image features with the cost volume and propagate the cost volume information to broader receptive fields. Specifically, the multi-scale encoder takes in $\bm{F}_{\mathrm{cv}}^{t}$ and the output is concatenated with $\{\bm{F}_{s}^{t}\}$ before sending into a UNet++ [48]-like decoder to upsample to full resolution and predict a depth candidates map $\bm{D}_{c}^{t}\in\mathbb{R}^{K\times H\times W}$, and Gaussian triplet map $\bm{F}^{t}_{l}\in\mathbb{R}^{C\times H\times W}$. We then predict the depth map through soft-argmax to bound the depth prediction between near and far:
$\bm{D}^{t}=\sum_{k=1}^{K}\mathrm{softmax}(\bm{D}_{c}^{t})_{k}\cdot d_{k}.$ (5)
Finally, the pixel-aligned Gaussian triplet map $\bm{F}^{t}_{l}$ is unprojected to 3D Gaussian triplet $\{\bm{\mu}_{l}^{t},\bm{\omega}_{l}^{t},\bm{f}^{t}_{l}\}$, where $\bm{\mu}_{l}^{t}\in\mathbb{R}^{3\times HW}$ are the Gaussian centers, $\bm{\omega}_{l}^{t}\in\mathbb{R}^{1\times HW}$ are weights between $(0,1)$, and $\bm{f}^{t}_{l}\in\mathbb{R}^{(C-1)\times HW}$ are Gaussian triplet features.
4.3 Pixel-wise Triplet Fusion
One limitation of previous generalizable 3DGS methods is the redundancy of Gaussians. Since we need multi-view observations to predict accurately localized 3D Gaussians in indoor scenes, the pixel-aligned Gaussians become redundant in frequently observed regions. Furthermore, previous methods integrate multi-view Gaussians of the same region simply through their opacities, leading to suboptimal performance due to lack of post aggregation (cf. Table 5). Consequently, inspired by previous methods [31, 40], we propose the Pixel-wise Triplet Fusion (PTF) module which can significantly remove redundant Gaussians in the overlapping regions and explicitly aggregate multi-view observation features in the latent space. We align the per-view local Gaussians with global ones using Pixel-wise Alignment to select the redundant 3D Gaussian Triplets, and progressively fuse the local Gaussians into the global ones.
Pixel-wise Alignment. Given the Gaussian triplets $\{\bm{\mu}^{t}_{l},\bm{f}_{l}^{t}\}_{t=1}^{T}$, we start from $t=1$ where the global Gaussians latent is empty. In $t$-th step, we first project the global Gaussian triplet centers $\bm{\mu}_{g}^{t-1}\in\mathbb{R}^{3\times M}$ onto the $t$-th view:
$\bm{\mathrm{p}}_{g}^{t}:=\{\bm{\mathrm{x}}_{g}^{t},\bm{\mathrm{y}}_{g}^{t},\bm{\mathrm{d}}_{g}^{t}\}=\bm{P}^{t}\bm{\mu}_{g}^{t-1},$ (6)
where $[\bm{\mathrm{x}}_{g}^{t},\bm{\mathrm{y}}_{g}^{t},\bm{\mathrm{d}}_{g}^{t}]\in\mathbb{R}^{3\times M}$ are the projected 2D coordinates and corresponding depths. We then correspond the local Gaussian triplets with the pixel-wise nearest projections within a threshold. Specifically, for the $i$-th local Gaussian with 2D coordinate $[\bm{\mathrm{x}}_{l}^{t}(i),\bm{\mathrm{y}}_{l}^{t}(i)]$ and depth $\rm{\mathrm{d}}_{l}^{t}(j)$, we first find its intra-pixel global projection set $\bm{\mathcal{S}}_{i}$:
$\bm{\mathcal{S}}_{i}^{t}:=\{j\mid[\bm{\mathrm{x}}_{g}^{t}(j)]=\bm{\mathrm{x}}_{l}^{t}(i),[\bm{\mathrm{y}}_{g}^{t}(j)]=\bm{\mathrm{y}}_{l}^{t}(i)\},$ (7)
[…]
where $\delta$ is a ratio threshold. We define the valid correspondence set as:
[…]
Gaussian Triplet Fusion. After the pixel-wise alignment, we remove the redundant 3D Gaussians through merging the validly aligned triplet pairs. Given a pair $(i,m_{i})\in\rm{\mathcal{F}}^{t}$, to restrict the 3D Gaussian centers to lie between the triplet pair, we weighted sum their center coordinates and sum their weights:
$\bm{\mu}^{t}_{g}(m_{i})=\frac{\bm{\omega}^{t}_{l}(i)\bm{\mu}^{t}_{l}(i)+\bm{\omega}^{t-1}_{g}(m_{i})\bm{\mu}^{t-1}_{g}(m_{i})}{\bm{\omega}^{t}_{l}(i)+\bm{\omega}^{t}_{g}(m_{i})},$ (10)
$\bm{\omega}^{t}_{g}(m_{i})=\bm{\omega}^{t}_{l}(i)+\bm{\omega}^{t-1}_{g}(m_{i}).$ (11)
To aggregate the local and global Gaussian latent features, we update the global Gaussian latent feature through a lightweight GRU network:
$\bm{f}_{g}^{t}(m_{i})=\mathrm{GRU}(\bm{f}_{l}^{t}(i),\bm{f}_{g}^{t-1}(m_{i}))$ (12)
Then the other local Gaussian triplets are appended to the global ones.
Gaussian primitives decoding. After the Pixel-wise Triplet Fusion, we can decode the global Gaussian triplets into Gaussian primitives:
$\bm{\Sigma},\bm{\alpha},\bm{\mathrm{s}}=\mathrm{MLP}_{d}(\bm{f}_{g}^{T})$ (13)
and the Gaussian centers $\bm{\mu}=\bm{\mu}_{g}^{T}$. Our proposed fusion method can incrementally integrate the Gaussians with geometrical constraints and learnable GRU network for feature update, which is capable of significantly removing redundant Gaussians and perform post feature aggregation across multiple views, and can be trained with the other framework components end-to-end with eligible computation overhead.
4.4 Training
Loss Functions. After predicting the 3D Gaussian primitives, we render from novel views following the rendering equations in Eq. (2). Similar to pixelSplat [1] and MVSplat [2], we train our framework using only photometric losses, i.e. a combination of MSE loss and LPIPS [49] loss, with weights of 1 and 0.05 following [1, 2].
Free-View Training. We propose a Free-View Training (FVT) strategy to add more geometrical constraints on the localization of 3D Gaussians, and to disentangle the performance of generalizable 3DGS with specific number of input views. To this end, we randomly sample $T$ number of context views (in experiments we set $T$ between $2$ and $8$), and supervise the image renderings in the broader view interpolations. The long sequence training is made feasible due to our efficient feature extraction and aggregation. We empirically find that FVT significantly contributes to depth estimation from novel views (cf. Table 3, 4).
5 Experiments
5.1 Experimental Settings
Datasets. We leverage the real-world indoor dataset ScanNet [18] for training. ScanNet is a large RGB-D dataset containing $1,513$ indoor scenes with camera poses, and we follow [50, 40] to use 100 scenes for training and 8 scenes for testing. To evaluate the generalization ability of our model, we further perform zero-shot evaluation on the synthetic indoor dataset Replica [19], for which we follow [51] to select 8 scenes for testing.
Implementation Details. Our FreeSplat is trained end-to-end using Adam [52] optimizer with an initial learning rate of $41e-$ 4 for 300k iterations following [1, 2]. Due to the large GPU requirements of [1, 2] given high-resolution images, all input images are resized to $512\times 384$ and batch size is set to 1, to form a fair comparison between different methods. We mainly compare with previous generalizable 3DGS methods in 2, 3, 10 reference view settings to evaluate the models’ performance under different length of input views. For 10 views setting, we also choose target views that are beyond the given sequence of reference views to evaluate the view extrapolation results. Furthermore, we notice that in MVSplat there is a misalignment between the depth plane candidates and the unprojection (more discussion in Appendix 5), which we modified and report their improved version in the following unless otherwise specified.
Table 1: Generalizable Novel View Interpolation results on ScanNet [18]. Except for FreeSplat-fv, other methods are all trained on specific number of views to form a complete comparison. Time(s) indicates the total time of encoding input images and rendering one image.
Method 2 views 3 views
[…]
NeuRay [9] 25.65 0.840 0.264 3.103 - 25.47 0.843 0.264 4.278 -
pixelSplat [1] 26.03 0.784 0.265 0.289 1180 25.76 0.782 0.270 0.272 1769
MVSplat [2] 25.05 0.771 0.258 0.105 393 24.24 0.769 0.273 0.192 590
FreeSplat-spec 27.70 0.826 0.222 0.095 292 27.24 0.820 0.232 0.101 384
FreeSplat-fv 27.18 0.818 0.226 0.097 296 26.79 0.813 0.238 0.104 393
Table 2: Long Sequence (10 views) Explicit Reconstruction results on ScanNet. The results of pixelSplat, MVSplat and FreeSplat-spec are given using their 3-views version.
Method Time(s)$\downarrow$ #GS(k) View Interpolation View Extrapolation
[…]
pixelSplat [1] 0.948 5898 21.26 0.714 0.396 20.70 0.687 0.429
MVSplat [2] 1.178 1966 18.87 0.664 0.417 18.72 0.642 0.444
FreeSplat-spec 0.273 1071 23.06 0.755 0.354 22.30 0.722 0.391
FreeSplat-fv 0.307 958 24.91 0.792 0.278 24.05 0.773 0.304
Table 3: Novel View Depth Rendering results on ScanNet. : 10-views results of pixelSplat, MVSplat and FreeSplat-spec are given using their 3-views version.
Method 2 views 3 views 10 views
[…]
NeuRay [9] 0.358 0.200 0.755 0.231 0.117 0.873 0.202 0.108 0.875
pixelSplat [1] 1.205 0.745 0.472 0.698 0.479 0.836 0.970 0.621 0.647
MVSplat [2] 0.526 0.274 0.396 0.394 0.201 0.625 0.423 0.240 0.558
FreeSplat-spec 0.239 0.121 0.833 0.218 0.108 0.882 0.214 0.122 0.850
FreeSplat-fv 0.220 0.117 0.847 0.209 0.099 0.888 0.159 0.092 0.917
5.2 Results on ScanNet
View Interpolation Results. On ScanNet, we evaluate the generalizable novel view interpolation results given 2 and 3 reference views as shown in Table 1. Comparing to pixelSplat and MVSplat, our FreeSplat-spec consistently improves PSNR by over $1.67$ dB on 2-views setting and $1.48$ dB on 3-views setting. Although slightly underperforming on SSIM comparing to NeuRay [9], we show significant improvements on PSNR and LPIPS over NeuRay and 300×300$\times$300 × faster inference speed. Moreover, our FreeSplat-fv consistently offers competitive results given arbitrary number of views.
Figure 3: Qualitative Results of Long Sequence Explicit Reconstruction. For each sequence, the first two rows are view interpolation results, and the last two rows are view extrapolation results.
Long Sequence Results. We further evaluate the long sequence results where we sample reference views with length of 10, and compare both view interpolation and extrapolation results as shown in Table 2. The results reveal that generalizable 3DGS methods underperform when given long sequence input images, which is due to the complicated camera trajectories in ScanNet, and the inaccuracy of 3D Gaussian localization that leads to errors when observed beyond the interpolated views. Our FreeSplat-3views still outperforms pixelSplat and MVSplat by over $1.80$ dB PSNR on view interpolation and $1.60$ dB PSNR on view extrapolation results. Through our proposed FVT that can be easily plugged into our model due to our low requirement on GPU, our FreeSplat-fv significantly outperforms our 3-views version by over $1.85$ dB PSNR. Our PTF module can also reduce the number of Gaussians by $45.5\%-51.3\%$, which significantly contributes to the efficiency of 3D Gaussians when given long sequence inputs. The qualitative results are shown in Figure 3, where the results clearly reveal that FreeSplat-spec outperforms MVSplat and pixelSplat in localizing 3D Gaussian and preserving fine-grained details, and FreeSplat-fv further improves on localizing and fusing multi-view Gaussians.
Novel View Depth Estimation Results. We also investigate the correctness of 3D Gaussian localization of different methods through comparing their depth rendering results. We report the Absolute Difference (Abs. Diff), Relative Difference (Rel. Diff), and threshold tolerance $\delta<1.25$ results from novel views in Table 3. We find that FreeSplat significantly outperforms pixelSplat and MVSplat in predicting accurately localized 3D Gaussians, where FreeSplat-fv reaches $91.7\%$ of $\delta<1.25$, outperforming pixelSplat and MVSplat by over $27.0\%$, enabling accurate unsupervised depth estimation from novel views. The improved depth estimation accuracy of FreeSplat-fv highlights the importance of depth estimation in supporting free-view synthesis across broader view range.
Table 4: Zero-Shot Transfer Results on Replica [19].
Method 3 Views 10 Views
[…]
pixelSplat [1] 26.24 0.829 0.229 0.576 1769 19.23 0.719 0.414 0.375 5898
MVSplat [2] 21.51 0.749 0.267 0.295 590 16.06 0.644 0.459 0.203 1966
FreeSplat-spec 27.08 0.848 0.175 0.657 415 19.83 0.735 0.361 0.585 1571
FreeSplat-fv 26.28 0.834 0.193 0.655 429 20.99 0.757 0.322 0.665 1478
Table 5: Ablation on ScanNet. CV: Cost Volume, PTF: Pixel-wise Triplet Fusion, FVT: Free-View Training.
CV PTF FVT 3 views 10 views
[…]
26.92 0.810 0.241 0.233 0.869 22.58 0.739 0.363 0.231 0.836
22.10 0.696 0.359 0.393 0.639 17.94 0.607 0.487 0.447 0.543
27.24 0.820 0.232 0.218 0.882 23.06 0.755 0.354 0.214 0.850
25.94 0.790 0.254 0.237 0.860 24.22 0.760 0.310 0.169 0.899
26.79 0.813 0.238 0.209 0.879 24.91 0.792 0.278 0.159 0.917
5.3 Zero-Shot Transfer Results on Replica
We further evaluate the zero-shot transfer results through testing on Replica dataset, with results in Table 4. Our view interpolation and novel view depth estimation results still outperforms existing methods. The long sequence results degrade due to inaccurate depth estimation and domain gap, indicating potential future work in further improving the depth estimation in zero-shot tranferring.
Figure 4: Qualtitative Ablation Study. The first and second row use input view lengths of 3 and 10.
5.4 Ablation Study
We conduct a detailed ablation study as shown in Table 5 and Figure 4. The results indicate that cost volume is essential in accurately localizing 3D Gaussians, and our proposed PTF module can consistently contribute to rendering quality and depth estimation results through its post aggregation of Gaussians. Furthermore, our FVT module excels in long sequence reconstruction results as well as novel view depth rendering results, indicating its significance in the unsupervised learning of depth estimation and reconstructing geometrically sound scenes.
6 Conclusion
In this study, we introduced FreeSplat, a generalizable 3DGS model that is tailored to accommodate an arbitrary number of input views and perform free-view synthesis using the global 3D Gaussians. We developed a Low-cost Cross-View Aggregation pipeline that enhances the model’s ability to efficiently process long input sequences, thus incorporating stricter geometry constraints. Additionally, we have devised a Pixel-wise Triplet Fusion module that effectively reduces redundant pixel-aligned 3D Gaussians in overlapping regions and merges multi-view Gaussian latent features. FreeSplat consistently improves the fidelity of novel view renderings in terms of both color image quality and depth map accuracy, facilitating feed-forward global Gaussians reconstruction without depth priors.
Appendix A Appendix / supplemental material
A.1 Experimental Environment
We conduct all the experiments on single NVIDIA RTX A6000 GPU. The experimental environment is PyTorch 2.1.2 and CUDA 12.2.
A.2 Additional Implementation Details
We set the number of virtual depth planes $K=64$, matching feature dimension $C_{m}=16$, and $d_{near}=0.5,d_{far}=15.0$ for cost volume formulation, and set $\delta=0.05$ in Eq.(8) for pixel-wise alignment. During our training and testing, the interval between reference views is set to 20 for ScanNet and 10 for Replica. To train the 3-views version of pixelSplat [1] on a single NVIDIA RTX A6000 GPU, we change their ViT patch size from $8\times 8$ to $16\times 16$. During inference on the 10 views setting, the epipolor line sampling in pixelSplat and the cross-view attention and cost volume formulation in MVSplat [2] are performed between nearby views similarly as ours to save GPU requirements and form a fair comparison, where for 3-views verion of pixelSplat, MVSplat and FreeSplat we set $N=2$ to match their training settings, and for the free-view version of FreeSplat we set $N=4$.
Table 6: Ablation on cost volume framework design.
3 views
[…]
MVSplat [2] 24.24 0.769 0.273 0.394 0.625
MVSplat [2] w/o Matched Depth Candidates 24.02 0.754 0.276 0.422 0.574
FreeSplat-spec w/o Matching Net (MN) 26.46 0.801 0.249 0.254 0.840
FreeSplat-spec w/o Multi-Scale Decoder (MSD) 26.07 0.799 0.281 0.263 0.825
FreeSplat-spec w/o (MN+MSD) 24.94 0.767 0.306 0.300 0.776
FreeSplat-spec 27.24 0.820 0.232 0.218 0.882
A.3 Additional Ablation Study
Figure 5: Mismatched depth plane candidates and unprojection in MVSplat [2].
Comparison with MVSplat. Since MVSplat [2] shares commons with our method in high-level design of building cost volume for generalizable 3DGS, we explore the key design differences that leads to our performance improvements. As shown in Figure 5, we notice that MVSplat uses the predicted depth as the ray distance instead of the distance to the image plane, which results in mismatching between the cost volume and point unprojection. As shown in Table 6, we evaluate the effectiveness of our matching net and the multi-scale decoder to compare with MVSplat. We also evaluate the influence of the mismatching problem on MVSplat. We use the image features $\bm{F}_{s=1}^{t}$ for cost volume formulation when the matching net is removed. We upsample and concatenate the cost volume with $\bm{F}_{s=0}^{t}$ to predict pixel-aligned Gaussian triplets when the multi-scale decoder is removed. The results indicate that removing the matching net (MN) and the multi-scale decoder (MSD) can result in decreased performance on novel view synthesis and depth estimation. Our performance becomes similar as MVSplat when both MN and MSD are removed, which indicates that our cost volume encoding and decoding is more suitable for aggregating more nearby views and can work more effectively on real-world image sequences with complicated camera poses.
Table 7: Ablation on the correlation between performance and the number of input views.
[…]
FreeSplat-3$\rightarrow$2views 25.70 0.808 0.240 0.293 0.747
FreeSplat-8$\rightarrow$2views 26.35 0.803 0.242 0.228 0.842
FreeSplat-2views 27.70 0.826 0.222 0.239 0.833
FreeSplat-fv 27.18 0.818 0.226 0.220 0.847
FreeSplat-8$\rightarrow$10views 24.92 0.788 0.276 0.153 0.926
FreeSplat-$\rightarrow$10views 24.91 0.792 0.278 0.159 0.917
Correlation between view number and performance. To evaluate the influence of changing the number of input views during inference on our method, we conduct an ablation study as shown in Table 7. The results indicate that training on one specific number of input views can lead to notiable performance decrease when inference on different number of views. However, our free-view training strategy can mitigate such correlation, such that to fit to arbitrary number of input views. On 10-views setting, our FreeSplat-fv can still perform similarly as FreeSplat-8views, indicating that using varying number of input views does not harm the long sequence reconstruction results.
A.4 Additional Qualitative Results
2 and 3-View Interpolation Results. The qualitative results are shown in Figure 6, where FreeSplat more precisely localizes 3D Gaussians and captures more fine-grained details comparing to previous methods. FreeSplat can also localize 3D Gaussians more accurately and renders precise depth maps, supporting high-quality rendering from broader view range (cf. FreeSplat-spec results in Figure 3).
Figure 6: Qualtitative Results given 2 and 3 reference views. We show the rendered color images (first row) and depth maps (second row) for each batch of reference views.
Results on Replica. We show the qualitative results on Replica in Figure 8, where our superiority over MVSplat and pixelSplat remains. The results indicate the generalization ability of FreeSplat across indoor datasets for the view interpolation task.
Figure 7: Qualitative Results on Replica.
Figure 8: Qualitative Results of whole scene reconstruction.
Results of Whole Scene Reconstruction. We also show qualitative results of our whole scene reconstruction in Figure 7. Despite the long input sequence ($\sim 40$ images) covering the whole scene, FreeSplat can still perform efficient feed-forward in $\sim 1\mathrm{s}$ on single NVIDIA RTX A6000, and can render high-quality images and accurate depth maps from novel views. On the other hand, it’s still difficult to accurately predict depth of textureless (e.g. wall) and specular (e.g. light reflection on the floor) regions. However, we hope our work provides an initial step towards accurate geometry reconstruction without ground truth depth priors.
A.5 Limitations
Although our approach excels in novel view rendering depth estimation and support arbitrary number of input views, the GPU requirement becomes expensive ($>40$GB) when inputting extremely long image sequence ($>50$). On the other hand, due to our unsupervised scheme of depth estimation, there is still a gap between our 3D reconstruction accuracy and the state-of-the-art methods with 3D supervision [33, 32] or RGB-D inputs [36, 35]. Our main focus is to explore the feed-forward indoor scene photorealistic reconstruction purely based on 2D supervision.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: We mainly focus on generalizable 3DGS for indoor scene reconstruction, where we support arbitrary number of inputs through adaptive cost volume (Sec. 4.2) and gaussian fusion (Sec. 4.3).
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: Discussed in Sec. A.5
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: We do not propose new theory in this paper.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We include detailed description of our framework in Sec. 4 and implementation details in Sec. 5.1.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [No]
  24. Justification: Our code will be released upon paper acceptance.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: The detailed descriptions of the experiment settings are illustrated in Sec. 5.1.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [No]
  34. Justification: We follow our related works in the setting for error bars.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We report the experimental environment in Sec. A.1.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: To the best of our knowledge, our work is conducted with NeurIPS Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [N/A]
  49. Justification: To the best of our knowledge, we do not foresee societal impacts of our work.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification: Our work does not pose such risks.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We properly cite the used existing datasets and models.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We provide detailed descriptions about our method (Sec. 4) as well as its limitations (Sec. A.5). Our code will be released upon paper acceptance.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: We do not involve crowdsourcing research.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: We do not involve crowdsourcing research.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.